DIRC Operations Report

Monday 2007/12/12 to Wednesday 2007/12/19


Oncall: Jose Benitez (Rolf Andreassen on backup)
Fast Monitoring: Rolf Andreassen, Jose Benitez

Contents:

    PEP Activities     BaBar Activities     DIRC activities     Backgrounds     Scheduled activities     Pending Work     Links

Luminosities this week

PEP peak lumi: 3.5 x 1033 (last week: 0 x 1034)

PEP delivery: .57 fb-1 (last week: 0 fb-1)

BaBar integrated lumi:.53 fb-1 (last week: 0 fb-1)

Run 7 total delivery: .57 fb-1

Run 7 total BaBar recorded: .53 fb-1 total, including 0 fb-1 off-peak data

PEP activities/issues this week

  • PEP has started giving us luminosity, although it is still very low.
  • They were mainly scrubbing and had some instabilities with the longitudinal feed back system and with a vacuum problem.
  • Scrubbing is done and now they are just trying to push the HER and LER currents up.

    BaBar activities/issues since last DIRC meeting

  • Some computing problems due to poor electrical power in the computer building otherwise ready to take data.
  • There was a power dip in IR2 on Monday, but there were no real consequences other than a few DAQ crates being turned off in the EH.

    DIRC activities/issues this week

  • On Thursday a new relay monitoring the power line in the Chiller was installed to try to remove the glitches. We are not sure that the problem is fixed because the ambient temperatures are not very cold anymore.
  • On Friday the Jerry made controlled access to the top of the detector to investigate the instabilities of the N2 flow meters. The instabilities go away as soon as he starts touching the electronics. Since then they have been stable.

    See DIRC logbook for details

    Lumi and Background this week